OpenStudy (anonymous):

What is black body radiation?? And its distribution of energy spectrum radiation..

OpenStudy (kropot72):

The information here will help your understanding: https://en.wikipedia.org/wiki/Black-body_radiation

OpenStudy (michele_laino):

the black body radiation is the radiation which is emitted by a "black body", and a "black body" is defined as the body whose absorbing power is equal to 1. the black body radiation is described by the Planck's law, the intensity of the black body radiation depends on the wave length of the emitted radiation and on the temperature of the body which emits such radiation

OpenStudy (anonymous):

Okay thanks..

OpenStudy (michele_laino):

the energy distribution of the black body radiation, as I wrote before, is given by the Planck's law: \[\Large E\left( {\lambda ,T} \right) = \left( {\frac{{8\pi hc}}{{{\lambda ^5}}}} \right)\frac{1}{{\exp \left( {\frac{{hc}}{{\lambda kT}}} \right) - 1}}\] where E is the energy density (energy over volume) over wave length. T is the absolute temperature and k is the Boltzamnn constant

OpenStudy (michele_laino):

h is the Planck's constant

OpenStudy (michele_laino):

Boltzmann*

OpenStudy (anonymous):

Umm okay now i get it.. and quantum theory of radiation is??

OpenStudy (michele_laino):

the quantum theory of radiation, is substantially the application of the quantum mechanics to light

OpenStudy (anonymous):

I am asking about max planks quantum theory of radiation..

OpenStudy (michele_laino):

I think that the Planck's theory is represented by his famous formula, since at his time quantum mechanics did not exist yet

OpenStudy (anonymous):

E is directly proportional to frequency.. is this formula for his law..??

OpenStudy (michele_laino):

yes! It is

OpenStudy (anonymous):

Okay thankssss..